THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 115 — 2 February 2012

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Figure 1: The Small Magellanic Cloud and Galactic globular cluster 47 Tucanae, two of the most beautiful objects in the Southern sky (from Kalirai et al. 2012; South is up and East is to the right). Image courtesy of and \begin{matrix}c \\ ◯ \end{matrix} Stephane Guisard, http://www.astrosurf.com/sguisard

Editorial

Dear Colleagues,

It is my pleasure to present you the 115^{th} issue of the Magellanic Clouds Newsletter. There is a lot of attention to star clusters, pulsating stars, binary stars, and the dynamics of the Magellanic System; don’t miss some of the very interesting proceedings papers.

It is not yet too late to register for the workshop on mass return from stars to galaxies, at STScI in March, but abstracts need to be in by 3 February (!): http://www.stsci.edu/institute/conference/mass-loss-return

The next meeting which should be of interest to readers of this newsletter, is the Magellanic Clouds meeting in Perth, Australia in September – see the announcement at the end of the newsletter for more details.

Thanks to those of you who supplied illustrations for this issue. They remain welcome. Also, don’t hesitate to use the ”announcement” type of submission to post results, ideas, requests for observations, et cetera. This is a forum, for discussion and collaboration.

The next issue is planned to be distributed on the 1^{st} of April 2012.

Editorially Yours, Jacco van Loon

Figure 2: Finderchart of the 30 Doradus (Tarantula Nebula) region, for stars of which spectra are shown in Figure 3.

Figure 3: Spectra of some rapidly-rotating stars with spectral classifications – see Walborn et al. (2012), on page 14. 3

Refereed Journal Papers

A new super-soft X-ray source in the Small Magellanic Cloud: Discovery of the first Be/white dwarf system in the SMC?

R. Sturm^{1}, F. Haberl^{1}, W. Pietsch^{1}, M.J. Coe^{2}, S. Mereghetti^{3}, N. La Palombara^{3}, R.A. Owen^{4} and A. Udalski^{5}

1Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße, 85748 Garching, Germany 2School of Physics and Astronomy, University of Southampton, Highfield, Southampton SO17 1BJ, United Kingdom 3INAF, Istituto di Astrofisica Spaziale e Fisica Cosmica Milano, via E. Bassini 15, 20133 Milano, Italy 4Laboratoire AIM, CEA/DSM – CNRS – Université Paris Diderot, IRFU/SAp, CEA-Saclay, 91191 Gif-sur-Yvette, France 5Warsaw University Observatory, Aleje Ujazdowskie 4, 00-478 Warsaw, Poland

The Small Magellanic Cloud (SMC) hosts a large number of Be/X-ray binaries, however no Be/white dwarf system is known so far, although population synthesis calculations predict that they might be more frequent than Be/neutron star systems. XMMU J010147.5−715550 was found as a new faint super-soft X-ray source (SSS) with a likely Be star optical counterpart. We investigate the nature of this system and search for further high-absorbed candidates in the SMC. We analysed the XMM-Newton X-ray spectrum and light curve, optical photometry, and the I-band OGLE III light curve. The X-ray spectrum is well represented by black-body and white dwarf atmosphere models with highly model-dependent temperature between 20 and 100 eV. The likely optical counterpart AzV 281 showed low near infrared emission during X-ray activity, followed by a brightening in the I-band afterwards. We find further candidates for high-absorbed SSSs with a blue star as counterpart. We discuss XMMU J010147.5−715550 as the first candidate for a Be/white dwarf binary system in the SMC.

Accepted for publication in A&A Available from arXiv:1112.0176

Metallicity distribution functions of the old populations of the Magellanic Clouds from RR Lyrae stars

Raoul Haschke^{1}, Eva K. Grebel^{1}, Sonia Duffau^{1} and Shoko Jin^{1}

1Astronomisches Rechen-Institut, Zentrum für Astronomie der Universität Heidelberg, Mönchhofstr. 12–14, D-69120 Heidelberg, Germany

We present the first metallicity distribution functions of the old field populations in the Magellanic Clouds. Our metallicities are based on the Fourier decomposition of Type ab RR Lyrae light curves from the Optical Gravitational Lensing Experiment (OGLE-III). On the metallicity scale of Zinn & West; we find a mean metallicity of [Fe/H] = −1.50 ± 0.24 dex based on 16776 RR Lyrae stars in the Large Magellanic Cloud (LMC). For the Small Magellanic Cloud (SMC) we obtain −1.70 ± 0.27 dex based on 1831 RR Lyrae stars. These uncertainties represent the intrinsic spread in the population rather than the standard deviation of the mean. Our results are in good agreement with the few existing spectroscopic metallicity determinations for LMC RR Lyrae stars from the literature. For both the LMC and the SMC the metallicity spread exceeds 1 dex in [Fe/H]. The distribution of metallicities in both Clouds is very uniform, and no significant metallicity gradient is detectable. We also do not find any pronounced populations of extremely metal-poor RR Lyrae candidates with metallicities well below −2 dex, although we need to caution that the photometric method used may overestimate the metallicities of metal-deficient stars. Moreover, because of stellar evolutionary effects one does not expect to observe many RR Lyrae stars among very metal-poor horizontal branch stars. We suggest that the Magellanic Clouds experienced fairly rapid and efficient early enrichment involving preenriched gas as well as possibly gas infall, while metal loss through outflows does not seem to have played a significant role. Moreover we suggest that the differences in the metallicities of the old population of LMC and SMC make an origin from a single, common progenitor unlikely, unless the separation happened very early on.

Accepted for publication in Astronomical Journal Available from arXiv:1111.6586

Globular cluster abundances from high-resolution, integrated-light spectroscopy. IV. The Large Magellanic Cloud: α, Fe-peak, light, and

heavy elements

J.E. Colucci^{1}, R.A. Bernstein^{1}, S.A. Cameron^{2} and A. McWilliam^{3}

1University of California Santa Cruz, USA 2Cerro Coso, USA 3Carnegie Observatories, USA

We present detailed chemical abundances in 8 clusters in the Large Magellanic Cloud (LMC). We measure abundances of 22 elements for clusters spanning a range in age of 0.05 to 12 Gyr, providing a comprehensive picture of the chemical enrichment and star formation history of the LMC. The abundances were obtained from individual absorption lines using a new method for analysis of high resolution (R ∼ 25, 000) integrated light spectra of star clusters. This method was developed and presented in Papers I, II, and III of this series. In this paper, we develop an additional integrated light χ^{2}-minimization spectral synthesis technique to facilitate measurement of weak (∼ 15 mÅ) spectral lines and abundances in low signal-to-noise ratio data (S/N ∼ 30). Additionally, we supplement the integrated light abundance measurements with detailed abundances that we measure for individual stars in the youngest clusters (Age < 2 Gyr) in our sample.In both the integrated light and stellar abundances we find evolution of [α/Fe] with [Fe/H] and age. Fe-peak abundance ratios are similar to those in the Milky Way, with the exception of [Cu/Fe] and [Mn/Fe], which are sub-solar at high metallicities. The heavy elements Ba, La, Nd, Sm, and Eu are significantly enhanced in the youngest clusters. Also, the heavy to light s-process ratio is elevated relative to the Milky Way ([Ba/Y] > +0.5) and increases with decreasing age, indicating a strong contribution of low-metallicity AGB star ejecta to the inter-stellar medium throughout the later history of the LMC. We also find a correlation of integrated light Na and Al abundances with cluster mass, in the sense that more massive, older clusters, are enriched in the light elements Na and Al with respect to Fe, which implies that these clusters harbor star-to-star abundance variations as is common in the Milky Way. Our results can be used to constrain both future chemical evolution models for the LMC and theories of GC formation.

Accepted for publication in Astrophysical Journal Available from arXiv:1111.6068

SXP 1062, a young Be X-ray binary pulsar with long spin period; Implications for the neutron star birth spin

F. Haberl^{1}, R. Sturm^{1}, M.D. Filipović^{2}, W. Pietsch^{1} and E.J. Crawford^{2}

1Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße, D-85748 Garching, Germany 2University of Western Sydney, Locked Bag 1797, Penrith South DC, NSW1797, Australia

The Small Magellanic Cloud (SMC) is ideally suited to investigating the recent star formation history from X-ray source population studies. It harbours a large number of Be/X-ray binaries (Be stars with an accreting neutron star as companion), and the supernova remnants can be easily resolved with imaging X-ray instruments. We search for new supernova remnants in the SMC and in particular for composite remnants with a central X-ray source. We study the morphology of newly found candidate supernova remnants using radio, optical and X-ray images and investigate their X-ray spectra. Here we report on the discovery of the new supernova remnant around the recently discovered Be/X-ray binary pulsar CXO J012745.97−733256.5 = SXP 1062 in radio and X-ray images. The Be/X-ray binary system is found near the centre of the supernova remnant, which is located at the outer edge of the eastern wing of the SMC. The remnant is oxygen-rich, indicating that it developed from a type Ib event. From XMM-Newton observations we find that the neutron star with a spin period of 1062 s (the second longest known in the SMC) shows a very high average spin-down rate of 0.26 s per day over the observing period of 18 days. From the currently accepted models, our estimated age of around 10000–25000 years for the supernova remnant is not long enough to spin down the neutron star from a few 10 ms to its current value. Assuming an upper limit of 25000 years for the age of the neutron star and the extreme case that the neutron star was spun down by the accretion torque that we have measured during the XMM-Newton observations since its birth, a lower limit of 0.5 s for the birth spin period is inferred. For more realistic, smaller long-term average accretion torques our results suggest that the neutron star was born with a correspondingly longer spin period. This implies that neutron stars in Be/X-ray binaries with long spin periods can be much younger than currently anticipated.

Accepted for publication in A&A Letters Available from arXiv:1112.0491

A deep, wide-field, and panchromatic view of 47 Tuc and the SMC with HST: Observations and data analysis methods

Jason S. Kalirai^{1,2}, Harvey B. Richer^{3}, Jay Anderson^{1}, Aaron Dotter^{1}, Gregory G. Fahlman^{4}, Brad M.S. Hansen^{5}, Jarrod Hurley^{6}, Ivan R. King^{7}, David Reitzel^{5}, R. Michael Rich^{5}, Michael M. Shara^{8}, Peter B. Stetson^{4} and Kristin

A. Woodley^{3}

1STScI, USA 2CAS, JHU, USA 3UBC, USA 4NRC/HIA, Canada 5UCLA, USA 6Swinburne, Australia 7UW, USA 8AMNH, USA

In Hubble Space Telescope (HST) Cycle 17, we imaged the well known globular star cluster 47 Tucanae for 121 orbits using the Wide Field Channel (WFC) of the Advanced Camera for Surveys (ACS) and both the UVIS and IR channels of the newly installed Wide Field Camera 3 (WFC3) instrument (GO-11677, PI – H. Richer). This unique data set was obtained by many scientific questions that demand a very deep, panchromatic, and panoramic view of the cluster’s stellar populations. In total, the program obtained over 0.75 Ms of imaging exposure time with the three HST cameras, over a time span of 9 months in 2010. The primary ACS field was imaged in the two broadband filters F 606W and F 814W filters, at 13 orientations, for all 121 orbits. The parallel WFC3 imaging provides a panchromatic (0.4–1.7 μm) and contiguous imaging swath over a 250^{∘} azimuthal range at impact radii of 6.5–17.9 pc in 47 Tuc. This imaging totals over 60 arcmin^{2} in area and utilizes the F 390W and F 606W broadband filters on WFC3/UVIS and the F 110W and F 160W broadband filters on WFC3/IR. In this paper, we describe the observational design of the new survey and one of the methods used to analyze all of the imaging data. This analysis combines over 700 full-frame images taken with the three HST cameras into a handful of ultra-deep, well-sampled combined images in each of the six filters. We discuss in detail the methods used to calculate accurate transformations that provide optimal alignment of the input images, the methods used to perform sky background offsets in the input stack and the flagging of deviant pixels, and the balance reached between the input pixel drop size onto an output super-sampled pixel grid. Careful photometric, morphological, and astrometric measurements are performed on the stacks using iterative PSF-fitting techniques, and reveal unprecedented colormagnitude diagrams (CMDs) of the cluster extending to > 30^{th} magnitude in the optical, 29^{th} magnitude in the UV, and 27^{th} magnitude in the IR. The data set provides a characterization of the complete stellar populations of 47 Tuc, extending from the faintest hydrogen burning dwarfs through the main-sequence and giant branches, down to very cool white dwarf remnants in the cluster. The imaging also provides the deepest probe of the stellar populations of the background Small Magellanic Cloud (SMC) galaxy, resolving low mass main-sequence dwarfs with M< ∼0.2 M_{⊙}.

Accepted for publication in Astronomical Journal, 143, 11 (2012) Available from arXiv:1112.1426

Principal Component Analysis on chemical abundances spaces

Y.S. Ting^{1,2}, K.C. Freeman^{1}, C. Kobayashi^{1,3}, G.M. De Silva^{4} and J. Bland-Hawthorn^{5}

1Research School of Astronomy & Astrophysics, The Australian National University, Cotter Road, Weston Creek, ACT 2611, Australia 2Department of Physics, National University of Singapore, Singapore 119260 3Centre for Astrophysics Research, University of Hertfordshire, Hatfield AL10 9AB, United Kingdom 4Australian Astronomical Observatory, P.O. Box 296, NSW 1710, Australia 5Institute of Astronomy, School of Physics, University of Sydney, NSW 2006, Australia

In preparation for the HERMES chemical tagging survey of about a million Galactic FGK stars, we estimate the number of independent dimensions of the space defined by the stellar chemical element abundances [X/Fe]. This leads to a way to study the origin of elements from observed chemical abundances using Principal Component Analysis. We explore abundances in several environments, including solar neighbourhood thin/thick disk stars, halo metalpoor stars, globular clusters, open clusters, the Large Magellanic Cloud and the Fornax dwarf spheroidal galaxy. By studying solar-neighbourhood stars, we confirm the universality of the r-process that tends to produce [n-capture elements/Fe] in a constant ratio. We find that, especially at low metallicity, the production of r-process elements is likely to be associated with the production of α-elements. This may support the core-collapse supernovae as the r-process site. We also verify the over-abundances of light s-process elements at low metallicity, and find that the relative contribution decreases at higher metallicity, which suggests that this lighter elements primary process may be associated with massive stars. We also verify the contribution from the s-process in low-mass AGB stars at high metallicity. Our analysis reveals two types of core-collapse supernovae: one produces mainly α-elements, the other produces both α-elements and Fe-peak elements with a large enhancement of heavy Fe-peak elements which may be the contribution from hypernovae. Excluding light elements that may be subject to internal mixing, K and Cu, we find that the [X/Fe] chemical abundance space in the solar neighbourhood has about 6 independent dimensions both at low metallicity (−3.5 < [Fe/H] < −2) and high metallicity ([Fe/H] > −1). However the dimensions come from very different origins in these two cases. The extra contribution from low-mass AGB stars at high metallicity compensates the dimension loss due to the homogenization of the core-collapse supernovae ejecta. Including the extra dimensions from [Fe/H], K, Cu and the light elements, the number of independent dimensions of the [X/Fe]+[Fe/H] chemical space in the solar neighbourhood for HERMES is about 8 to 9. Comparing fainter galaxies and the solar neighbourhood, we find that the chemical space for fainter galaxies such as Fornax and the Large Magellanic Cloud has a higher dimensionality. This is consistent with the slower star formation history of fainter galaxies. We find that open clusters have more chemical space dimensions than the nearby metal-rich field stars. This suggests that a survey of stars in a larger Galactic volume than the solar neighbourhood may show about 1 more dimension in its chemical abundance space.

Accepted for publication in Monthly Notices of the Royal Astronomical Society Available from arXiv:1112.3207

The Large Magellanic Cloud and the distance scale

Alistair R. Walker^{1}

1Cerro Tololo Inter-American Observatory, National Optical Astronomy Observatory, Chile

The Magellanic Clouds, especially the Large Magellanic Cloud, are places where multiple distance indicators can be compared with each other in a straight-forward manner at considerable precision. We here review the distances derived from Cepheids, Red Variables, RR Lyraes, Red Clump Stars and Eclipsing Binaries, and show that the results from these distance indicators generally agree to within their errors, and the distance modulus to the Large Magellanic Cloud appears to be defined to ±3% with a mean value (m − M)_{0} = 18.48 mag, corresponding to 49.7 Kpc. The utility of the Magellanic Clouds in constructing and testing the distance scale will remain as we move into the era of Gaia.

Accepted for publication in Astrophysics and Space Science Available from arXiv:1112:3171

Evolution and hydrodynamics of the very-broad X-ray line emission in SN 1987A

D. Dewey^{1}, V.V. Dwarkadas^{2}, F. Haberl^{3}, R. Sturm^{3} and C.R. Canizares^{1}

1MIT Kavli Institute, USA 2University of Chicago, USA 3MPE Garching, Germany

Deep Chandra and XMM-Newton observations show that the interaction of SN 1987A with its environment gives rise to X-ray emission from multiple components: the H ii region, the dense protrusions of the equatorial ring (ER) including clumpy ER material, and, increasingly at later times, the reverse-shocked supernova (SN) ejecta itself. The High Energy Transmission Grating (HETG) observation in 1999 showed very broad lines with a full-width at halfmaximum (FWHM) of order 7000 km s^{−1}; at this time the blast wave was primarily interacting with the H ii region around the progenitor. Since then, the X-ray emission has been increasingly dominated by narrower components as the blast wave encounters the dense ER. Even so, a very-broad emission line component in the grating spectra suggests that flux from the H ii region is still present. We find that deep HETG 2007 and 2011 data are better fit when of order 20% of the flux comes from a component with FWHM around 9000 km s^{−1}. Based on this, the observed X-ray light curves in the 0.5–2 and 3–10 keV bands are modeled as the weighted sum of the non-equilibrium-ionization (NEI) X-ray emission from two simple 1D hydrodynamic simulations: one for the H ii region interaction producing very-broad emission lines, and one for the dense equatorial ring collision. Using these simple models we interpret the X-ray emission components and their properties (e.g., abundances), predict future light curve changes under different model assumptions, and set initial parameters for a multi-dimensional model of SN 1987A’s interaction.

Submitted to Astrophysical Journal Available from arXiv:1111.5314

XSL: The X-Shooter Spectral Library

Yanping Chen^{1}, Scott Trager ^{1}, Reynier Peletier^{1} and Ariane Lançon^{2}

1Kapteyn Astronomical Institute, University of Groningen, Postbus 800, 9700 AV, Groningen, Netherlands 2Observatoire Astronomique, 11, rue de l’Université, F-67000 Strasbourg, France

We are building a new spectral library with the X-Shooter instrument on ESO’s VLT: XSL, the X-Shooter Spectral Library. We present our progress in building XSL, which covers the wavelength range from the near-UV to the near-IR with a resolution of R ∼ 10, 000. At the time of writing we have collected spectra for nearly 240 stars. An important feature of XSL is that we have already collected spectra of more than 100 Asymptotic Giant Branch stars in the Galaxy and the Magellanic Clouds.

Published in Journal of Physics: Conference Series Available from arXiv:1112.3651

The role of dwarf galaxy interactions in shaping the Magellanic System and implications for Magellanic Irregulars

Gurtina Besla^{1,2}, Nitya Kallivayalil^{3}, Lars Hernquist^{2}, Roeland P. van der Marel^{4}, T.J. Cox^{5} and Dušan Kereš^{6}

1Columbia University, USA 2Harvard CfA, USA 3Yale University, USA 4STScI, USA 5Carnegie Observatories, USA 6UC Berkeley, USA

We present a novel pair of numerical models of the interaction history between the Large and Small Magellanic Clouds (LMC and SMC, respectively) and our Milky Way (MW) in light of recent high precision proper motions from the Hubble Space Telescope (Kallivayalil et al. 2006ab). These proper motions imply that the Magellanic Clouds (MCs) are moving ∼ 80 km s^{−1} faster than previously considered. Given these velocities, cosmological simulations of hierarchical structure formation favor a scenario where the MCs are currently on their first infall towards our Galaxy (Besla et al. 2007; Boylan-Kolchin et al. 2011; Busha et al. 2011). We illustrate here that the observed irregular morphology and internal kinematics of the Magellanic System (in gas and stars) are naturally explained by interactions between the LMC and SMC, rather than gravitational interactions with the MW. These conclusions provide further support that the MCs are completing their first infall to our system. In particular, we demonstrate that the Magellanic Stream, a band of H i gas trailing behind the Clouds 150^{∘} across the sky, can be accounted for by the action of LMC tides on the SMC before the system was accreted by the MW. We further demonstrate that the off-center, warped stellar bar of the LMC and its one-armed spiral, can be naturally explained by a recent direct collision with its lower mass companion, the SMC. Such structures are key morphological characteristics of a class of galaxies referred to as Magellanic Irregulars (de Vaucouleurs & Freeman 1972), the majority of which are not associated with massive spiral galaxies. We infer that dwarf–dwarf galaxy interactions are important drivers for the morphological evolution of Magellanic Irregulars and can dramatically affect the efficiency of baryon removal from dwarf galaxies via the formation of extended tidal bridges and tails. Such interactions are important not only for the evolution of dwarf galaxies but also have direct consequences for the buildup of baryons in our own MW, as LMC-mass systems are believed to be the dominant building blocks of MW-type halos.

Accepted for publication in MNRAS Available from arXiv:1201.1299

Eccentric ellipsoidal red-giant binaries in the LMC: Complete orbital solutions and comments on interaction at periastron

C.P. Nicholls^{1} and P.R. Wood^{2}

1Center for Astrophysics and Space Science, University of California San Diego, La Jolla CA 92093, USA 2Research School for Astronomy and Astrophysics, Australian National University, Cotter Road, Weston Creek ACT 2611 Australia

Modelling ellipsoidal variables with known distances can lead to exact determination of the masses of both components, even in the absence of eclipses. We present such modelling using light and radial velocity curves of ellipsoidal red giant binaries in the LMC, where they are also known as sequence E stars. Stars were selected as likely eccentric systems on the basis of light curve shape alone. We have confirmed their eccentric nature and obtained system parameters using the Wilson–Devinney code. Most stars in our sample exhibit unequal light maxima as well as minima, a phenomenon not observed in sequence E variables with circular orbits. We find evidence that the shape of the red giant changes throughout the orbit due to the high eccentricity and the varying influence of the companion. Brief intervals of pulsation are apparent in two of the red giants. We determine pulsation modes and comment on their placement in the period–luminosity plane. Defining the parameters of these systems paves the way for modelling to determine by what mechanism eccentricity is maintained in evolved binaries.

Accepted for publication in MNRAS Available from arXiv:1201.1043

The clustered nature of star formation. Pre–main-sequence clusters in the star-forming region NGC 602/N 90 in the Small Magellanic Cloud

Dimitrios A. Gouliermis^{1,2}, Stefan Schmeja^{3}, Andrew E. Dolphin^{4}, Mario Gennaro^{1}, Emanuele Tognelli^{5,6} and Pier Giorgio Prada Moroni^{5,6}

1Max Planck Institute for Astronomy, Königstuhl 17, 69117 Heidelberg, Germany 2Zentrum für Astronomie der Universität Heidelberg, Inst. für Theoretische Astrophysik, Albert-Überle-Str. 2, 69120 Heidelberg, Germany 3Zentrum für Astronomie der Universität Heidelberg, Astronomisches Rechen-Institut, Mönchhofstr. 12–14, 69120 Heidelberg, Germany 4Raytheon Company, PO Box 11337, Tucson, AZ 85734, USA 5Dipartimento di Fisica ”Enrico Fermi”, Università di Pisa, largo Pontecorvo 3, Pisa I-56127, Italy 6INFN-Sezione di Pisa, largo Pontecorvo 3, Pisa I-56127, Italy

Located at the tip of the wing of the Small Magellanic Cloud (SMC), the star-forming region NGC 602/N 90 is characterized by the H ii nebular ring N 90 and the young cluster of pre–main-sequence (PMS) and early-type main sequence stars NGC 602. We present a thorough cluster analysis of the stellar sample identified with HST/ACS camera in the region. We show that apart from the central cluster, low-mass PMS stars are congregated in thirteen additional small compact sub-clusters at the periphery of NGC 602. We find that the spatial distribution of the PMS stars is bimodal, with an unusually large fraction (∼ 60%) of the total population being clustered, while the remaining is diffusely distributed in the inter-cluster area. From the corresponding color–magnitude diagrams we disentangle an age-difference of ∼ 2.5 Myr between NGC 602 and the compact sub-clusters which appear younger. The diffuse PMS population appears to host stars as old as those in NGC 602. Almost all detected PMS sub-clusters appear to be centrally concentrated. When the complete PMS stellar sample, including both clustered and diffused stars, is considered in our cluster analysis, it appears as a single centrally concentrated stellar agglomeration, covering the whole central area of the region. Considering also the hot massive stars of the system, we find evidence that this agglomeration is hierarchically structured. Based on our findings we propose a scenario, according to which the region NGC 602/N 90 experiences an active clustered star formation for the last ∼ 5 Myr. The central cluster NGC 602 was formed first and rapidly started dissolving into its immediate ambient environment, possibly ejecting also massive stars found away from its center. Star formation continued in sub-clusters of a larger stellar agglomeration, introducing an age-spread of the order of 2.5 Myr among the PMS populations.

Accepted for publication in The Astrophysical Journal Available from arXiv:1201.3081

Blood ties: the real nature of the LMC binary globular clusters NGC 2136 and NGC 2137

A. Mucciarelli ^{1}, L. Origlia^{2}, F.R. Ferraro^{1}, M. Bellazzini^{2} and B. Lanzoni^{1}

1Department of Astronomy, Bologna University, Italy 2INAF – Osservatorio Astronomico di Bologna, Italy

We have used a sample of high-resolution spectra obtained with the multi-fiber facility FLAMES at the Very Large Telescope of the European Southern Observatory, to derive the kinematical and chemical properties of the two young Large Magellanic Cloud globular clusters NGC 2136 and NGC 2137. These two clusters represent a typical example of LMC cluster pair suspected to be bound in a binary system: indeed the cluster centers of gravity have an angular separation of < 1.4′ in the sky. The spectral analysis of seven giants in NGC 2136 and four in NGC 2137 reveals that the two clusters share very similar systemic radial velocities, namely v_{rad} = 271.5 ± 0.4 km s^{−1} (σ = 1.0 km s^{−1}) and v_{rad} = 270.6 ± 0.5 km s^{−1} (σ = 0.9 km s^{−1}) for NGC 2136 and NGC 2137, respectively, and they have also indistinguishable abundance patterns. The iron content is [Fe/H] = −0.40 ± 0.01 dex (σ = 0.03 dex) for NGC 2136 and −0.39 ± 0.01 dex (σ = 0.01 dex) for NGC 2137, while the [α/Fe] ratios are roughly solar in both clusters. These findings suggest that the two clusters are gravitationally bound and that they formed from the fragmentation of the same molecular cloud that was chemically homogeneous. This is the first firm confirmation of the binary nature of a LMC cluster pair. The most likely fate of this system is to merge into a single structure in a time-scale comparable with its orbital period.

Accepted for publication in ApJ Letters Available from arXiv:1201.3715

A global physical model for Cepheids

Ondřej Pejcha^{1} and Christopher S. Kochanek^{1,2}

1Department of Astronomy, The Ohio State University, 140 West 18^{th} Avenue, Columbus, OH 43210, USA 2Center for Cosmology and Astroparticle Physics, The Ohio State University, 191 West Woodruff Avenue, Columbus, OH 43210, USA

We perform a global fit to ∼ 5, 000 radial velocity and ∼ 177, 000 magnitude measurements in 29 photometric bands covering 0.3 μm to 8.0 μm distributed among 287 Galactic, LMC, and SMC Cepheids with P > 10 days. We assume that the Cepheid light curves and radial velocities are fully characterized by distance, reddening, and time-dependent radius and temperature variations. We construct phase curves of radius and temperature for periods between 10 and 100 days, which yield light curve templates for all our photometric bands and can be easily generalized to any additional band. With only 4 to 6 parameters per Cepheid, depending on the existence of velocity data and the amount of freedom in the distance, the models have typical rms light and velocity curve residuals of 0.05 mag and 3.5 km s^{−1}. The model derives the mean Cepheid spectral energy distribution and its derivative with respect to temperature, which deviate from a black body in agreement with metal-line and molecular opacity effects. We determine a mean reddening law towards the Cepheids in our sample, which is not consistent with standard assumptions in either the optical or near-IR. Based on stellar atmosphere models we predict the biases in distance, reddening, and temperature determinations due to the metallicity and we quantify the metallicity signature expected for our fit residuals. The observed residuals as a function of wavelength show clear differences between the individual galaxies, which are compatible with these predictions. In particular, we find that metal-poor Cepheids appear hotter. Finally, we provide a framework for optimally selecting filters that yield the smallest overall errors in Cepheid parameter determination, or filter combinations for suppressing or enhancing the metallicity effects on distance determinations. Light curve templates and a code to find the best fit period and reference time as well as accurate mean magnitudes are available at http://www.astronomy.ohio-state.edu/∼pejcha/cepheids. The code uses our light curve templates and hence mean magnitudes are calculated even for bands with a single datapoint.

Accepted for publication in ApJ Available from arXiv:1112.3038 and from http://www.astronomy.ohio-state.edu/∼pejcha/cepheids

CORS Baade–Wesselink distance to the LMC NGC 1866 blue populous cluster

Roberto Molinaro^{1}, Vincenzo Ripepi^{1}, Marcella Marconi^{1}, Ilaria Musella^{1}, Enzo Brocato^{2}, Alessio Mucciarelli^{3}, Peter B. Stetson^{4}, Jesper Storm^{5} and Alistair Walker^{6}

1INAF – Osservatorio Astronomico Capodimonte, Napoli, Italy 2INAF – Osservatorio Astronomico di Collurania, Teramo, Italy 3Dipartimento di Astronomia, Bologna, Italy 4Dominion Astrophysical Observatory, Herzberg Institute of Astrophysics, National Research Council, Victoria, Canada 5Astrophysikalisches Institut, Potsdam, Germany 6Cerro Tololo Inter-American Observatory, National Optical Astronomy Observatory, La Serena, Chile

We used optical, near-infrared photometry and radial velocity data for a sample of 11 Cepheids belonging to the young LMC blue populous cluster NGC 1866 to estimate their radii and distances on the basis of the CORS Baade–Wesselink method. This technique, based on an accurate calibration of the surface brightness as a function of (U − B), (V − K) colors, allows us to estimate, simultaneously, the linear radius and the angular diameter of Cepheid variables, and consequently to derive their distance. A rigorous error estimate on radius and distances was derived by using Monte Carlo simulations. Our analysis gives a distance modulus for NGC 1866 of 18.51 ± 0.03 mag, which is in agreement with several independent results.

Accepted for publication in Astrophysical Journal Available from arXiv:1201.3478

The dust budget of the SMC: Are AGB stars the primary dust source at low metallicity?

Martha L. Boyer^{1}, S. Srinivasan^{2}, D. Riebel^{3}, I. McDonald^{4}, J.Th. van Loon^{5}, G.C. Clayton^{6}, K.D. Gordon^{1}, M. Meixner^{1}, B.A. Sargent^{1} and G.C. Sloan^{7}

1Space Telescope Science Institute, USA 2Institut d’Astrophysique de Paris, France 3Department of Physics and Astronomy, The Johns Hopkins University, USA 4Jodrell Bank Centre for Astrophysics, University of Manchester, UK 5Astrophysics Group, Lennard-Jones Laboratories, Keele University, UK 6Louisiana State University, Department of Physics & Astronomy, USA 7Astronomy Department, Cornell University, USA

We estimate the total dust input from the cool evolved stars in the Small Magellanic Cloud (SMC), using the 8 μm excess emission as a proxy for the dust-production rate. We find that Asymptotic Giant Branch (AGB) and red supergiant (RSG) stars produce (8.6–9.5)×10^{7} M_{⊙} yr^{−1} of dust, depending on the fraction of far-infrared sources that belong to the evolved star population (with 10%–50% uncertainty in individual dust-production rates). RSGs contribute the least (< 4%), while carbon-rich AGB stars (especially the so-called ”extreme” AGB stars) account for 87%–89% of the total dust input from cool evolved stars. We also estimate the dust input from hot stars and supernovae (SNe), and find that if SNe produce 10^{−3} M_{⊙} of dust each, then the total SN dust input and AGB input are roughly equivalent. We consider several scenarios of SNe dust production and destruction and find that the interstellar medium (ISM) dust can be accounted for solely by stellar sources if all SNe produce dust in the quantities seen around the dustiest examples and if most SNe explode in dense regions where much of the ISM dust is shielded from the shocks. We find that AGB stars contribute only 2.1% of the ISM dust. Without a net positive contribution from SNe to the dust budget, this suggests that dust must grow in the ISM or be formed by another unknown mechanism.

Accepted for publication in ApJ Available from arXiv:1201.5384

An X-ray study of supernova remnant N 49 and soft γ-ray repeater 0526−66 in the Large Magellanic Cloud

Sangwook Park^{1}, John P. Hughes^{2}, Patrick O. Slane^{3}, David N. Burrows^{4}, Jae-Joon Lee^{5} and Koji Mori^{6}

1University of Texas at Arlington, USA 2Rutgers University, USA 3Harvard-Smithsonian Center for Astrophysics, USA 4Pennsylvania State University, USA 5Korea Astronomy and Space Science Institute, South Korea 6University of Miyazaki, Japan

We report on the results from our deep Chandra observation (120 ks) of the supernova remnant (SNR) N 49 and soft γ-ray repeater (SGR) 0526−66 in the Large Magellanic Cloud. We firmly establish the detection of an ejecta ”bullet” beyond the southwestern boundary of N 49. The X-ray spectrum of the bullet is distinguished from that of the main SNR shell, showing significantly enhanced Si and S abundances. We also detect an ejecta feature in the eastern shell, which shows metal overabundances similar to those of the bullet. If N 49 was produced by a core-collapse explosion of a massive star, the detected Si-rich ejecta may represent explosive O-burning or incomplete Si-burning products from deep interior of the supernova. On the other hand, the observed Si/S abundance ratio in the ejecta may favor Type Ia origin for N 49. We refine the Sedov age of N 49, τ_{Sed} ∼ 4800 yr, with the explosion energy E_{0} ∼ 1.8 × 10^{51} erg. Our blackbody (BB) + power law (PL) model for the quiescent X-ray emission from SGR 0526−66 indicates that the PL photon index (Γ ∼ 2.5) is identical to that of PSR 1E 1048.1−5937, the well-known candidate transition object between anomalous X-ray pulsars and SGRs. Alternatively, the two-component BB model implies X-ray emission from a small (R ∼ 1 km) hot spot(s) (kT ∼ 1 keV) in addition to emission from the neutron star’s cooler surface (R ∼ 10 km, kT ∼ 0.4 keV). There is a considerable discrepancy in the estimated column toward 0526−66 between BB+PL and BB+BB model fits. Discriminating these spectral models would be crucial to test the long-debated physical association between N 49 and 0526−66.

Accepted for publication in ApJ Available from arXiv:1201.5056

UV extinction towards a quiescent molecular cloud in the SMC

J. Maíz Apellániz^{1} and M. Rubio^{2}

1IAA-CSIC, Granada, Spain 2Universidad de Chile, Santiago, Chile

Context: The mean UV extinction law for the Small Magellanic Cloud (SMC) is usually taken as a template for lowmetallicity galaxies. However, its current derivation is based on only five stars, thus placing doubts on its universality. An increase in the number of targets with measured extinction laws in the SMC is necessary to determine its possible dependence on parameters such as metallicity and star-forming activity. Aims: To measure the UV extinction law for several stars in the quiescent molecular cloud SMC B1-1. Methods: We obtained HST/STIS slitless UV spectroscopy of a 25′′ × 25′′ field of view and we combined it with ground-based NIR and visible photometry of the stars in the field. The results were processed using the Bayesian photometric package chorizos to derive the visible-NIR extinction values for each star. The unextinguished Spectral Energy Distributions (SEDs) obtained in this way were then used to derive the UV extinction law for the four most extinguished stars. We also recalculated the visible–NIR extinction for the five SMC stars with preexisting UV extinction laws. Results: The UV extinction law for four SMC B1-1 stars within several pc of each other differs significantly from star to star. The 2175 Å bump is moderately strong in one, weak in two, and absent in the fourth.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1201.6230

Photometric and spectroscopic studies of massive binaries in the Large Magellanic Cloud. I. Introduction and orbits for two detached systems:

Evidence for a mass discrepancy?

Philip Massey^{1}, Nidia I. Morrell^{2}, Kathryn F. Neugent^{1}, Laura R. Penny^{3}, Kathleen DeGioia-Eastwood^{4} and Douglas R. Gies^{5}

1Lowell Observatory, USA 2Las Campanas Observatory, Carnegie Observatories, USA 3Department of Physics and Astronomy, The College of Charleston, USA 4Department of Physics and Astronomy, Northern Arizona University, USA 5Center for High Angular Resolution Astronomy and Department of Physics and Astronomy, Georgia State University, USA

The stellar mass–luminosity relation is poorly constrained by observations for high mass stars. We describe our program to find eclipsing massive binaries in the Magellanic Clouds using photometry of regions rich in massive stars, and our spectroscopic follow-up to obtain radial velocities and orbits. Our photometric campaign identified 48 early-type periodic variables, of which only 15 (31%) were found as part of the microlensing surveys. Spectroscopy is now complete for 17 of these systems, and in this paper we present analysis of the first two, LMC 172231 and ST 2-28, simple detached systems of late-type O dwarfs of relatively modest masses. Our orbit analysis yields very precise masses (about 2%), and we use tomography to separate the components and determine effective temperatures by model fitting, necessary for determining accurate (0.05–0.07 dex) bolometric luminosities in combination with the light-curve analysis. Our approach allows more precise comparisons with evolutionary theory than previously possible.

To our considerable surprise, we find a small, but significant, systematic discrepancy: all of the stars are slightly undermassive, by typically 11% (or over-luminous by 0.2 dex) compared to that predicted by the evolutionary models. We examine our approach for systematic problems, but find no satisfactory explanation. The discrepancy is in the same sense as the long-discussed and elusive discrepancy between the masses measured from stellar atmosphere analysis with the stellar evolutionary models, and might suggest that either increased rotation or convective overshooting is needed in the models. Additional systems will be discussed in future papers of this series, and will hopefully confirm or refute this trend.

Accepted for publication in Astrophysical Journal Available from arXiv:1201.3280

Conference Papers

Two remarkable spectroscopic categories of young O stars from the VLT-FLAMES Tarantula Survey

Nolan R. Walborn^{1}, Hugues Sana^{2}, William D. Taylor^{3}, Sergio Simón-Díaz^{4} and Christopher J. Evans^{5}

1Space Telescope Science Institute, USA 2Astronomical Institute Anton Pannekoek, University of Amsterdam, The Netherlands 3Scottish Universities Physics Alliance, Institute for Astronomy, University of Edinburgh, Royal Observatory Edinburgh, Scotland 4Instituto de Astrofísica de Canarias, Spain 5UK Astronomy Technology Centre, Royal Observatory Edinburgh, Scotland

The spectral and spatial characteristics of two special categories of O stars found in the VFTS dataset are presented. One of them comprises very rapid rotators, including several more extreme than any previously known. These objects are distributed around the peripheries of the main 30 Doradus clusters, suggesting a runaway nature for which their radial velocities already provide preliminary supporting evidence. The other category consists of a large number of Vz stars, previously hypothesized on spectroscopic grounds to be on or very near the ZAMS. Their distribution is the inverse of that of the rapid rotators: the Vz are strongly concentrated to the ionizing clusters, plus a newly recognized band of recent and current star formation to the north, which provides strong circumstantial evidence for their extreme youth.

Oral contribution, published in ”Four Decades of Research on Massive Stars”, eds. Laurent Drissen, Nicole St-Louis, Carmelle Robert & Anthony F.J. Moffat, ASPCS Available from arXiv:1112.4435

H.E.S.S. observations of the Large Magellanic Cloud

Nukri Komin^{1}, A. Djannati-Atai^{2}, Y. Gallant^{3}, V. Marandon^{2}, C.C. Lu^{4}, S. Ohm^{4,5,6} and E. de Ona-Wilhelmi^{4}

1LAPP, CNRS/IN2P3, Annecy-le-Vieux, France 2APC, CNRS/IN2P3, Paris, France 3LUPM, CNRS/IN2P3, Montpellier, France 4Max-Planck-Institut für Kernphysik, Heidelberg, Germany 5School of Physics & Astronomy, University of Leeds, UK 6Department of Physics & Astronomy, University of Leicester, UK

The Large Magellanic Cloud (LMC) is a satellite galaxy of the Milky Way at a distance of approximately 48 kpc. Despite its distance it harbours several interesting targets for TeV γ-ray observations. The composite supernova remnant N 157B/PSR J05367−6910 was discovered by H.E.S.S. being an emitter of very high energy (VHE) γ-rays.

It is the most distant pulsar wind nebula ever detected in VHE γ-rays. Another very exciting target is SN 1987A, the remnant of the most recent supernova explosion that occurred in the neighbourhood of the Milky Way. Models for Cosmic Ray acceleration in this remnant predict γ-ray emission at a level detectable by H.E.S.S. but this has not been detected so far. Fermi/LAT discovered diffuse high energy (HE) γ-ray emission from the general direction of the massive star forming region 30 Doradus but no clear evidence for emission from either N 157B or SN 1987A has been published. The part of the LMC containing these objects has been observed regularly with the H.E.S.S. telescopes since 2003. With deep observations carried out in 2010 a very good exposure of this part of the sky has been obtained. The current status of the H.E.S.S. LMC observations is reported along with new results on N 157B and SN 1987A.

Oral contribution, published in ”32^{nd} International Cosmic Ray Conference”, Beijing 2011 Available from arXiv:1201.0639

Microlensing towards the Magellanic Clouds and M 31: is the quest for MACHOs still open?

Sebastiano Calchi Novati^{1,2}

1Department of Physics, University of Salerno, Italy 2IIASS, Italy

Microlensing is the tool of choice for the search and the analysis of compact halo objects (“MACHOs”), a still viable class of dark matter candidates at the galactic scale. Different analyses point towards an agreement in excluding dark matter MACHOs of less than about 0.1 M_{⊙}; it remains however an ongoing debate for values in the mass range (0.1–1) M_{⊙}. The more robust constraints, though not all in agreement, come from the observational campaigns towards the Magellanic Clouds (the LMC and the SMC). The analyses towards the nearby galaxy of M 31, in the so called “pixel lensing” regime, have expanded the perspectives in this field of research. In this contribution first we draw a critical view on recent results and then we focus on the pixel lensing analysis towards M 31 of the PLAN collaboration.

Oral contribution, published in ”3^{rd} Italian–Pakistani Workshop on Relativistic Astrophysics” (Univ. of Lecce, Italy, June 2011), Journal of Physics: Conference Series (JPCS) Available from arXiv:1201.5262

SOUTH POL: Revealing the polarized southern sky

A.M. Magalhães^{1}, C.M. de Oliveira^{1}, A. Carciofi^{1}, R. Costa^{1}, E.M.G. Dal Pino^{1}, M. Diaz^{1}, T. Ferrari^{1}, C. Fernandez^{1}, A.L. Gomes^{1}, L. Marrara^{1}, A. Pereyra^{3}, N.L. Ribeiro^{1}, C.V. Rodrigues^{2}, M.S. Rubinho^{1}, D.B. Seriacopi^{1} and K. Taylor^{1}

1Instituto de Astronomia, Geofísica e Ciéncias Atmosféricas, Universidade de São Paulo, Rua do Matão 1226, São Paulo, SP 05508-090, Brazil 2Instituto Nacional de Pesquisas Espaciais/MCT, Av. dos Astronautas 1758, São José dos Campos, SP 12227-010, Brazil 3Observatório Nacional, Rua Gal. José Cristino 77, Rio de Janeiro, RJ 20921-400, Brazil

SOUTH POL will be a survey of the southern sky in optical polarized light. It will use a newly designed polarimetric module at an 80-cm robotic telescope. Telescope and polarimeter will be installed at CTIO, Chile, in late 2012. The initial goal is to cover the sky south of declination −15^{∘} in two years of observing time, aiming at a polarimetric accuracy < ∼0.1% down to V = 15 mag, with a camera covering a field of about 2.0 square degrees. SOUTH POL will impact areas such as cosmology, extragalactic astronomy, interstellar medium of the Galaxy and Magellanic Clouds, star formation, stellar envelopes, stellar explosions and Solar System, among others.

Poster contribution, published in ”Stellar Polarimetry: from Birth to Death”, eds. J. Hoffman, B. Whitney & J. Bjorkman, AIP Conf. Proc. (2012) Available from arXiv:1201.3639 and from http://www.aip.org (available after publication)

Announcement

Workshop on the Magellanic Clouds

Dear colleagues, We are pleased to announce a workshop on the subject of the Magellanic Clouds to be held in Perth from Sept. 10 to 13, 2012. The proposed scientific topics at the workshop include:

• The Magellanic Clouds in a cosmological context

• Proper motions of the Clouds and their past orbits

• Formation of the Magellanic System (simulation)

• Structure and formation of the Magellanic System (observation)

• Star formation histories and chemical properties

• Star formation and ISM

• From GMCs to stars and star clusters

• Structure and kinematics / stellar population The following link provides the scientific rationale and further information relating to the workshop: http://www.icrar.org/news/magellanic-clouds-workshop The registration will be open in February 2012, at which time we will also invite proposals for contributed talks and posters. We hope to see you here in September! Best Regards, Bi-Qing For (on behalf of the SOC) biqing.for@icrar.org See also http://www.icrar.org/news/magellanic-clouds-workshop